# Upgrade of the online data quality monitoring system for future PandaX experiments

### v1

@@ 1 Introduction
The PandaX-4T experiment [1] is a multi-tonne liquid xenon observatory located in the China Jinping Underground Laboratory  [2]. It utilizes the dual-phase liquid xenon technology [3] to explore a wide range of physics phenomena, such as dark matter detection [4], coherent elastic scattering of solar neutrinos [5, 6], double beta decay of xenon isotopes [7] and other rare events [8]. The future plans include the development of PandaX-xT, a multi-ten-tonne observatory that will operate at the same site [9].
Scattering events occurring in the active region of the dual-phase xenon detector may produce prompt scintillation light signal ($S1$) in liquid xenon and electroluminescence signal ($S2$) in gas xenon. These signals can be captured by two arrays of photomultiplier tubes (PMTs) on the top and bottom of the detector, read out separately with 169 and 199 channels. The captured signals are digitized using flash analog-to-digital converters (FADCs) and subsequently stored for further analysis. Due to the combination of high digitization rate at 250 MHz (500 MHz after upgrade [10] between Run1 and Run2 data taking process) and a large number of readout channels, a substantial amount of data is generated. It is estimated that PandaX-4T is collecting approximately 1 PB of data per year [11], while PandaX-xT is projected to produce over 1.8 PB of data per year [9]. These data are stored in a custom format based on the Bamboo Shoot3 library [12]. During the data taking process, an Apache Kafka message queue [13, 14, 15] is employed to transmit metadata information to various functional modules. These modules encompass the online data processing, data transfering and data quality monitoring system.
@@ 2 The data quality monitoring system
This section provides an introduction to the data quality monitoring system, which consists of three sub-modules: the plots generator, the app server and the web-based user interface (WebUI). The architecture of the system will be discussed, highlighting the roles and functionalities of each sub-module.
@@ 2.1 Architecture of the system
The data quality monitoring system is tightly integrated with the online data processing system and the database, as illustrated in Fig. 1. The interactions between its three sub-modules are distinctly represented.
@@
The different systems within the architecture are interconnected using the Apache Kafka message queue. This message queue allows for the creation and sharing of different “topics”, which serve as categories for organizing and sharing messages among the various systems involved [13, 14, 15]. In the DAQ system, raw data files are produced. After completing the writing process for a data file, a message containing the run and file number is sent to the “daq” topic. The online data processing system consumes this message and determines whether or not to process the raw data based on predefined rules. The data processing steps employed by the system follow the standard procedures introduced in Ref. [11]. Once the processing of a given file is completed, the same message is sent to the “dq” topic to notify the plots generator about extracting the data quality information. In the previous version of the plots generator, about 50 image files were generated for different variables that are valuable for quality monitoring, and some information is stored directly to the database.
@@ 2.2 Plots generator
For each selected raw data file, the online data processing system generates three output files, containing hit, signal, and event level information, respectively. Once the processing of a data file is completed, the plots generator is notified through the message queue. The multi-threaded plots generator operates with one thread dedicated to receiving messages and placing them into an internal data structure, which is a mapping between run numbers and the list of corresponding file numbers. Upon the arrival of a message for a new run, a dedicated thread is assigned to generate data quality plots for that particular run. The thread analyzes the output files in the order specified in the corresponding file number lists. Once all the related files have been analyzed, the file number is removed from the list, and the files are subsequently removed from storage. The thread is recycled for a new run when the file list remains empty for a minimum of one hour.
@@
In the previous version of the plots generator, the distributions were stored in memory using TH1D or TH2D objects from the ROOT [19] framework. These objects were updated as the input files were analyzed. The plots were generated in PNG format after processing every 10 raw files. The calculated variables were simultaneously uploaded to the database.
@@ 3 Upgrade of the plots generator
This section presents the updated version of the plots generator, which offers a significant increase in throughput compared to the initial version. The updated plots generator also includes support for the prompt acquisition of self-corrected PMT gain by utilizing online generated hit data.
@@ 5 Summary
Regarding the plots generator, we have made several updates. Firstly, we have revamped the internal data structure by replacing the histogram classes from ROOT with standard containers. Secondly, we have implemented multi-threading processing of incoming files, incorporating a locking mechanism to maintain data integrity. Additionally, we have transitioned from using image files to using JSON files for extracting data quality information. These updates have significantly improved the overall data throughput, allowing us to monitor half of the collected data online in the future operations of the PandaX-4T experiment.

### v2

# Upgrade of the PandaX-4T online data quality monitoring system and perspectives for future multi-tons PandaX upgrades

@@ 1 Introduction
The PandaX-4T experiment [1] is a multi-tonne liquid xenon observatory located in the China Jinping Underground Laboratory  [2]. It utilizes the dual-phase liquid xenon technology [3] to explore a wide range of physics phenomena, such as dark matter detection [4], coherent elastic scattering of solar neutrinos [5, 6], double beta decay of xenon isotopes [7] and other rare events [8]. After completing the initial data taking stage (Run1), a detector upgrade was carried out. The recent successful completion of the detector upgrade has paved the way for the scheduling of new data taking stage (Run2). The future plans include the development of PandaX-xT, a multi-ten-tonne observatory that will operate at the same site [9].
Scattering events occurring in the active region of the dual-phase xenon detector may produce prompt scintillation light signal ($S1$) in liquid xenon and electroluminescence signal ($S2$) in gas xenon. These signals can be captured by two arrays of photomultiplier tubes (PMTs) on the top and bottom of the detector, read out separately with 169 and 199 channels. The captured signals are digitized using flash analog-to-digital converters (FADCs) and subsequently stored for further analysis. Due to the combination of high digitization rate of 250 MHz (500 MHz post-upgrade [10]) and a large number of readout channels, a substantial amount of data is generated. It is estimated that PandaX-4T is collecting approximately 1 PB of data per year [11], while PandaX-xT is projected to produce over 1.8 PB of data per year [9]. These data are stored in a custom format based on the Bamboo Shoot3 library [12]. During the data taking process, an Apache Kafka message queue [13, 14, 15] is employed to transmit metadata information to various functional modules. These modules encompass the online data processing, data transfering and data quality monitoring system.
@@ 2 The data quality monitoring system
In this section, we introduce the architecture, components, and functionalities of the data quality monitoring system.
@@ 2.1 Architecture of the system
The data quality monitoring system is essential for swiftly identifying issues during data collection. It must extract physical information from processed data obtained through the online data processing system once the data is ready. Furthermore, the metadata of runs is stored in the database. Thus, the data quality monitoring system is tightly integrated with the online data processing system and the database, as illustrated in Fig. 1. The system comprises three independent sub-modules: a) the plots generator, responsible for extracting information from processed files; b) the Web UI, which serves as the user interface; c) the app server, reading extracted information from the database or file system and transmitting it to the user. The interactions among these sub-modules are distinctly represented in Fig. 1.
@@
The different systems within the architecture are interconnected using the Apache Kafka message queue. This message queue allows for the creation and sharing of different “topics”, which serve as categories for organizing and sharing messages among the various systems involved [13, 14, 15]. In the DAQ system, raw data from a single run are stored as a series of files, each assigned a unique file number for identification. After completing the writing process for a data file, a message containing the run and file number is sent to the “daq” topic. The online data processing system consumes this message and determines whether or not to process the raw data based on predefined rules. The data processing steps employed by the system follow the standard procedures introduced in Ref. [11]. Once the processing of a given file is completed, the same message is sent to the “dq”(data quality) topic to notify the plots generator about extracting the data quality information. In the previous version of the plots generator, about 50 image files were generated for different variables that are valuable for quality monitoring, and some information is stored directly to the database.
@@ 2.2 Plots generator
For each selected raw data file, the online data processing system generates three output files, containing hit, signal, and event level information, respectively. A hit corresponds to a pulse in the raw waveform of a PMT, containing information about detected photons. A signal is a grouping of hits from multiple PMTs with similar timing. An event consists of signals occurring within a specific time window, potentially originating from the same physical interaction [11]. Once the processing of a data file is completed, the plots generator is notified through the message queue. The multi-threaded plots generator operates with one thread dedicated to receiving messages and placing them into an internal data structure, which is a mapping between run numbers and the list of corresponding file numbers. Upon the arrival of a message for a new run, a dedicated thread is assigned to generate data quality plots for that particular run. The thread analyzes the output files in the order specified in the corresponding file number lists. Once all the related files have been analyzed, the file number is removed from the list, and the files are subsequently removed from storage. The thread is recycled for a new run when the file list remains empty for a minimum of one hour.
@@
In the previous version of the plots generator, the distributions were stored in memory using TH1D or TH2D objects from the ROOT [19] framework. These objects were updated as the input files were analyzed. The plots were regenerated in PNG format with updated information after processing every 10 raw files. The calculated variables were simultaneously uploaded to the database.
@@ 3 Upgrade of the plots generator
The updated plots generator offers a significant increase in throughput compared to the initial version. The updated plots generator also includes support for the prompt acquisition of self-corrected PMT gain by utilizing online generated hit data.
@@ 5 Summary
Regarding the plots generator, we have made several updates. Firstly, we have revamped the internal data structure by replacing the histogram classes from ROOT with standard containers. Secondly, we have implemented multi-threading processing of incoming files, incorporating a locking mechanism to maintain data integrity during concurrent updating of the run level information. Additionally, we have transitioned from using image files to using JSON files for extracting data quality information. These updates have significantly improved the overall data throughput, allowing us to monitor half of the collected data online in the future operations of the PandaX-4T experiment.